### v2

@@ Abstract
We introduce OneCAT, a unified multimodal model that seamlessly integrates understanding, generation, and editing within a novel, pure decoder-only transformer architecture. Our framework uniquely eliminates the need for external components such as Vision Transformers (ViT) or vision tokenizer during inference, leading to significant efficiency gains, especially for high-resolution inputs. This is achieved through a modality-specific Mixture-of-Experts (MoE) structure trained with a single autoregressive (AR) objective, which also natively supports dynamic resolutions. Furthermore, we pioneer a multi-scale visual autoregressive mechanism within the Large Language Model (LLM) that drastically reduces decoding steps compared to diffusion-based methods while maintaining state-of-the-art performance. Our findings demonstrate the powerful potential of pure autoregressive modeling as a sufficient and elegant foundation for unified multimodal intelligence. As a result, OneCAT sets a new performance standard, outperforming existing open-source unified multimodal models across benchmarks for multimodal generation, editing, and understanding.
@@ 1 Introduction
In the past few years, modular approaches, utilizing separate architectures for understanding [2, 69, 3, 13, 14, 12], generation [42, 37, 61, 24, 34, 20], and editing tasks [90, 5, 32, 92, 48, 91], dominated multi-modal frameworks. While producing capable systems, it inherently creates complex, multi-stage pipelines. Such designs often suffer from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance. In response to these limitations, the field has seen a rapid convergence towards unified multimodal modeling, aiming to integrate these disparate capabilities within a single, end-to-end architecture [71, 65, 80, 81, 11, 16]. Despite the trend towards unification, many current models remain tethered to the modular paradigm. We contend that a true paradigm shift—necessary to unlock the full potential of unified systems—demands a move towards a more fundamental, first-principles architecture that eschews external components. We therefore propose a pure decoder-only transformer, trained under a unified objective, providing not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
@@
To realize this vision, we re-evaluate the core architectural tenets of multimodal systems. We introduce an encoder-free framework where raw visual inputs are directly tokenized into patch embeddings and processed alongside text tokens within a single decoder stack. The critical innovation is a modality-specific Mixture-of-Experts (MoE) layer, which dynamically routes vision and text tokens to specialized experts, enabling deep, early-stage feature fusion in an efficient inference manner without the need for exquisite encoders. For generative tasks, we pioneer a multi-scale autoregressive mechanism [24, 66] inside the LLM, augmented with the proposed scale-aware adapter modules, to predict image tokens from low to high resolution. This design not only circumvents the high latency of iterative diffusion models but also allows the model to learn a coarse-to-fine generative process, significantly enhancing both speed and output quality.
Regarding data protocol, we leverage mixed training strategies that combine large-scale, web-scraped image-text pairs with curated, instruction-following datasets. Our pre-training corpus is constructed to expose the model to a diverse range of data formats, including image-text documents, visual question-answering pairs, and generative prompts for both text-to-image synthesis and image editing. This heterogeneous data mixture is crucial for training a truly unified model, as it forces the shared decoder to develop a generalized representation that can seamlessly switch between comprehension, generation, and editing tasks. We argue that this unified data diet, coupled with our bottleneck-free architecture, is key to fostering the synergistic emergence of complex multimodal capabilities.
Building upon these principles, we present Only DeCoder Auto-regressive Transformer (OneCAT), an open-source unified multimodal model. Comprehensive evaluations demonstrate that OneCAT sets a new state-of-the-art for pure decoder-only unified models. More importantly, the encoder-free design provides a significant inference speedup, particularly for high-resolution inputs. By demonstrating the viability and superiority of a pure decoder architecture, OneCAT offers a more first-principles-aligned paradigm for multimodal modeling. It facilitates earlier cross-modal fusion through its unified MoE structure and enhances semantic consistency, providing valuable insights and a powerful new baseline for the development of next-generation unified multimodal systems. For examples of our model’s impressive image generation capabilities, please refer to Fig. 1. Its advanced image editing functionalities are further showcased in Fig. 2.
@@ 2.1 Compositional MLLMs
The field of Multimodal Large Language Models (MLLMs) has rapidly evolved, converging on a dominant compositional architecture. This paradigm connects a pre-trained vision encoder, such as CLIP [58], SigLIP [89], or the more recent InternViT [14], to a powerful Large Language Model (LLM) through a trainable connector. Pioneering works [1, 39] propose sophisticated connector designs. For example, Flamingo [1] introduces gated cross-attention layers to inject visual information into an LLM, while BLIP-2 [39] develops the Q-Former to bridge the modality gap between an image encoder and an LLM. A significant shift occurs with LLaVA [47], which simplifies the connector to a lightweight Multi-Layer Perceptron (MLP) projection layer. This effective design become a foundational blueprint for numerous subsequent MLLMs. For example, recent state-of-the-art models like the InternVL series [14, 12, 13, 96, 70] and the Qwen-VL series [69, 3, 2] adopt this same core principle and achieve superior performance by leveraging larger-scale training data and more powerful vision and language foundation models. However, this successful compositional design has inherent drawbacks. The separate nature of the vision and language components complicates the end-to-end optimization process and introduces two critical bottlenecks. First, the sequential nature of the architecture, where the vision encoder must fully process an image before the LLM can begin its generation, leads to high inference latency. This initial step is often referred to as the “prefilling” stage. Second, the connector itself acts as an information bottleneck. In this so-called late fusion pipeline, complex visual information is compressed into a compact representation for the LLM, inevitably causing a loss of fine-grained visual detail. These fundamental limitations are now motivating a shift in the field towards more deeply integrated, such as decoder-only models, that aim to overcome these challenges.
@@ 2.2 Decoder-only MLLMs
Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the conventional compositional architecture. This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by eliminating the separate vision encoder. For example, Fuyu-8B [4] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM backbone, which markedly reduce inference latency. Inspired by this success, subsequent works [18, 51, 19, 35, 68, 62], further advance decode-only MLLMs by targeting their training processes and architectures. EvE [18] and VoRA [68] aligns the LLM’s hidden states with semantic features from a pre-trained ViT. However, directly using a smaller model (e.g., a ViT with hundreds of millions of parameters) as the teacher to distill knowledge into a significantly larger LLM (with several billion parameters) may restrict the expressive capacity of the LLM. Differently, Mono-InternVL [51] and EvEv2.0 [19] adopt a Mixture-of-Experts (MoE) framework, introducing a dedicated visual expert to handle visual-specific features more effectively. HoLVE [64] prepends a causal transformer to the LLM to explicitly convert both visual and textual inputs into a shared space. Despite these promising advancements, the overall training efficiency of these models remains a significant challenge. More importantly, the potential for the decoder-only architecture to create unified models that can seamlessly integrate multimodal understanding, generation, and even image editing capabilities remains a largely unexplored research avenue.
@@ 2.3 Unified Visual Understanding and Generation
Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier [73, 11, 79, 95, 71, 80, 38, 55, 8, 45, 16, 43]. Pioneering unified MLLMs such as Chameleon [65],Transfusion [95], emu3 [71], show-o [80] and Synergen-vl [38] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents. Janus series [73, 11] decouples visual encoding for understanding and generation using two separate encoders, but may compromise performance due to task conflicts in shared LLM parameter space. Metaqueries [55], BLIP3-O [8], Uniworld-V [45] assembles off-the-shelf specialized MLMMs and diffusion models by tuning adapters and learnable query tokens, which sacrifices true architectural unification for modularity. BAGEL [16] and Mogao [43] employ a Mixture-of-Transformers (MoT) architecture, dedicating different components to autoregressive text generation and diffusion-based visual generation. However, while powerful, this hybrid approach inherits the significant inference latency of diffusion models and still requires separate encoders and tokenizers during the inference.
@@ 2.4 Next Scale Prediction for Visual Generation
In contrast, our proposed OneCAT seamlessly unifies next-token prediction for text generation and next-scale prediction for visual generation within a single LLM.
@@ 3 OneCAT
Figure 3: Overview inference pipeline of OneCAT. OneCAT is a decoder-only autoregressive unified model that seamlessly supports multimodal understanding, text-to-image generation and image editing. OneCAT implements both next-token prediction for text generation and next-scale prediction for visual generation within a unified LLM backbone. In the left part, the dash squares denote the generated text and visual token during inference. In the right part, PE and $T$ denote patch embedding and text tokenizer, respectively. TD and ID denote text detokenizer and image detokenizer, respectively.
As illustrated in Fig. 3, OneCAT employs a pure decoder-only architecture, eliminating the need for any additional vision encoder or tokenizer during inference. This streamlined design significantly simplifies the model structure and reduces computational overhead. Unlike traditional multimodal models that rely on semantic encoders like ViTs for understanding [3, 16, 43, 11], OneCAT utilizes a lightweight Patch Embedding layer. This layer losslessly converts raw images into continuous visual tokens, enabling efficient multimodal understanding. Crucially, the same Patch Embedding layer also encodes reference images for editing tasks, thereby superseding separate VAE-based tokenizers traditionally used [16, 43] and further enhancing inference efficiency.
At its core, OneCAT integrates a Mixture-of-Experts (MoE) architecture. This MoE comprises three specialized feed-forward network (FFN) experts: one dedicated to processing text tokens for language comprehension, another designed for continuous visual tokens to facilitate multimodal understanding, and a third optimized for generating discrete visual tokens during image synthesis. All QKV and attention layers are shared across modalities and tasks, which promotes significant parameter efficiency and robust cross-modal alignment for instruction-following.
For text generation, OneCAT adheres to the conventional Next-Token Prediction paradigm, leveraging the well-established capabilities of autoregressive language modeling. In parallel, for visual generation, it innovatively employs the Next-Scale Prediction paradigm [66, 24]. This mechanism generates images in a coarse-to-fine, hierarchical manner, progressively predicting visual tokens from the lowest to the highest resolution scale, thereby achieving high-quality visual outputs.
3.1 Architecture
Our OneCAT model is initialized from the pre-trained Qwen2.5 LLM [82], leveraging its strong language modeling capabilities as a foundation. To construct our MoE architecture, we replicate the original FFN layer from each Qwen2.5 transformer block to form three distinct experts: a Text FFN (i.e., Text. FFN), a Visual Understanding FFN (i.e., Und. FFN), and a Visual Generation FFN (i.e., Gen. FFN). OneCAT employs a deterministic hard routing mechanism, where tokens are assigned to a specific expert based on their modality and the task at hand. The core functionality for each task is handled as follows:
@@ 3.1 Architecture
To process images for understanding tasks, we employ a simple yet effective patch embedding layer that converts raw images into a sequence of continuous visual tokens. This layer consists of a $\times14$ convolution for image patchifying, a $\times2$ pixel unshuffle operation for visual token compression, and a two-layer Multilayer Perceptron (MLP) for projecting the visual features to match the LLM’s hidden state dimension. These continuous visual tokens are exclusively routed to the Und. FFN, while the text tokens for instruction are routed to the Text FFN.
@@
For image generation, we leverage a pre-trained multi-scale VAE model from Infinity [24] to map images between pixel space and latent space. This VAE operates with a downsampling ratio of 16 and a latent channel size of 32, and incorporates a bitwise quantizer [94] to enlarge the vocabulary. During training, the image tokenizer transforms the target images into a sequence of multi-scale discrete visual tokens to serve as ground-truth and input of LLM for teacher-forcing training, which are processed by the Gen. FFN. Critically, during inference, the tokenizer is not required; only the detokenizer is needed to reconstruct the final image from the generated multi-scale visual tokens. The conditional text tokens are also routed to the Text FFN.
@@
OneCAT seamlessly supports image editing task by conditioning the visual generation process on a reference image and edit-instruction. The reference image is processed through the patch embedding layer, and the resulting continuous visual tokens are routed to the Und. FFN to serve as the visual condition, while the text tokens for instruction are routed to the Text FFN. The patch embedding layer provides a near-lossless representation of the reference image, which allows the LLM’s shallower layers to access low-level features for pixel-level consistency, while the deeper layers extract high-level features for semantical comprehension. Guided by this rich, hierarchical visual context, the model then autoregressively predicts new discrete visual tokens, which are handled by the Gen. FFN. This design enables powerful conditional generation without requiring any architectural modifications, showcasing the versatility of our unified decoder-only design.
@@ 3.2 Scale-Aware Adapter for Hierarchical Generation
While the Gen. FFN is capable of processing discrete visual tokens, the tokens produced by the multi-scale VAE are inherently hierarchical. A standard FFN would treat tokens for different scales equally, ignoring this crucial structural information. To address this and enable more granular control over the visual generation process, we introduce the Scale-Aware Adapter (SAA), a novel architectural component integrated with the Gen. FFN.
Our design is motivated by the principle that different scales in the multi-scale VAE govern distinct aspects for image generation. Specifically, tokens from lower scales predominantly encode low-frequency global information, such as color, illumination, and coarse structure. Those from higher scales, conversely, capture high-frequency details including fine textures and intricate patterns. Processing these functionally divergent tokens with a shared Gen. FFN layer limits representational capacity and is thus suboptimal.
The Scale-Aware Adapter (SAA) comprises a set of parallel modules that serve as skip connections over the Gen. FFN. Each module is dedicated to processing tokens from a specific scale of the multi-scale VAE, with the total number of adapters matching the number of VAE scales. During inference, discrete visual tokens are routed to their corresponding scale-specific adapter based on the scale index. To ensure parameter efficiency, each adapter is constructed using a low-rank decomposition (rank rr= 64), inspired by the LoRA [26]. However, unlike LoRA which is typically used for fine-tuning, the SAA modules are trained jointly end-to-end as permanent components of the LLM.

### v3

@@ Abstract
We introduce OneCAT, a unified multimodal model that seamlessly integrates understanding, generation, and editing within a single decoder-only transformer architecture. OneCAT uniquely eliminates the need for external components such as Vision Transformers (ViT) or vision tokenizer during inference, leading to significant efficiency gains, especially for high-resolution image inputs and outputs. This is achieved through a modality-specific Mixture-of-Experts (MoE) design trained with a unified autoregressive (AR) objective, which also natively supports dynamic resolutions. Furthermore, we pioneer to achieve multi-scale visual autoregressive mechanism within the Large Language Model (LLM) with proposed scale-aware adapter (SAA) that drastically reduces decoding latency compared to diffusion-based methods while maintaining state-of-the-art performance. Our findings demonstrate the powerful potential of pure autoregressive modeling as a sufficient and elegant foundation for unified multimodal intelligence. As a result, OneCAT outperforms existing unified models across benchmarks for multimodal understanding, generation, and editing.
[…]
@@ 1 Introduction
Modular approaches using separate modules for understanding [92, 5, 17], generation [81, 45, 25], and editing tasks [114, 7, 115] are becoming dominant for multi-modal frameworks. Despite producing capable systems, such designs inherently create complex multi-stage pipelines suffering from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance. Unified multimodal LLMs aim to address these limitations by integrating these disparate abilities within a single end-to-end architecture [94, 15, 21], but many methods remain tethered to the modular paradigm [73, 11]. This motivates us to achieve a true revolution on the fundamental architecture that unlocks the full potential of unified systems and eschews heavy external components. In this paper, we propose a single decoder-only autoregressive model trained under a unified objective to provide elegant and potent foundation for general-purpose multimodal intelligence.
Unified architecture. We propose the first encoder-free framework for unified multimodal LLM (MLLM), where raw visual inputs are directly tokenized into patch embeddings and are processed alongside text tokens within a single decoder model. The critical innovation is a modality-specific MoE layer that dynamically routes continuous vision tokens, discrete vision tokens, and text tokens to specialized experts. It enables deep and early-stage feature fusion without requiring exquisite encoders (i.e., ViT and vision tokenizer) for efficient inference. For generative tasks, we pioneer to embed the multi-scale autoregressive mechanism [89] into the LLM and propose a scale-aware adapter (SAA) to extract scale-specific representation for augmentation. Thus, image tokens can be predicted from low to high resolutions, while next-token prediction is adopted for text tokens. This design significantly enhances the speed and quality of image generation by simultaneously circumventing the high latency of diffusion models and learning a coarse-to-fine generative process.
@@
Unified training paradigm. Training encoder-free MLLMs for strong visual perception ability is notoriously data-intensive [64]. To mitigate this, existing methods like VoRA [91] and EvE [23] perform distillation to align the internal hidden states of an LLM student with a pre-trained ViT teacher. However, they suffer from supervisory bottleneck problem that the expressive capacity of an LLM is restricted by a small teacher. We address this problem with a novel distillation strategy that first customizes a powerful MLLM teacher and then efficiently transfers its comprehensive visual perception to the proposed encoder-free unified MLLM. On such basis, we use large-scale heterogeneous multimodal data and employ a unified expert pretraining, mid-training, and supervised fine-tuning to force the shared decoder to achieve generalized representation that can seamlessly switch between comprehension, generation, and editing tasks.
Building upon these innovations, we present Only DeCoder Auto-regressive Transformer (OneCAT), a unified multimodal model. Comprehensive evaluations demonstrate that OneCAT sets a new state-of-the-art for unified models. More importantly, the encoder-free design provides a significant inference speedup, particularly for high-resolution inputs and outputs. OneCAT demonstrates the viability and superiority of a pure decoder architecture, and offers a more first-principle-aligned paradigm for multimodal modeling. It facilitates earlier cross-modal fusion through its unified MoE structure and enhances semantic consistency via unified AR objective to provide valuable insights and a powerful new baseline for developments of next-generation unified multimodal systems. For examples of our model’s impressive image generation capabilities, please refer to Fig. 1. Its advanced image editing functionalities are further showcased in Fig. 2.
@@ 2.1 Compositional MLLMs
The field of Multimodal Large Language Models (MLLMs) has rapidly evolved, converging on a dominant compositional architecture. This paradigm connects a pre-trained vision encoder (i.e., CLIP [77], SigLIP [113], and InternViT [19]), to a powerful LLM through a trainable connector. Pioneering works [3, 51] propose sophisticated connector designs. For example, Flamingo [3] introduces gated cross-attention layers to inject visual information into an LLM, while BLIP-2 [51] develops the Q-Former to bridge the modality gap between the vision encoder and LLM. A significant shift occurs with LLaVA [60], which simplifies the connector to a lightweight MLP, which become a foundational blueprint for subsequent MLLMs. For example, recent state-of-the-art models like the InternVL series [19, 17, 18, 120, 93] and the Qwen-VL series [92, 5, 4] adopt this core principle and achieve superior performance by leveraging larger-scale training data and more powerful vision and language foundation models. However, this successful compositional design has inherent drawbacks. The separate nature of the vision and language components complicates the end-to-end optimization process and introduces two critical bottlenecks. First, the sequential nature of the architecture, where the vision encoder must fully process an image before the LLM can begin its generation, leads to high inference latency, especially for the prefilling stage. Second, the connector acts as an information bottleneck. In this so-called late fusion pipeline, complex visual information is compressed into a compact representation for the LLM, inevitably causing a loss of fine-grained visual detail. These fundamental limitations are now motivating a shift in the field towards more deeply integrated, such as decoder-only models, that aim to overcome these challenges.
@@ 2.2 Decoder-only MLLMs
Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the compositional MLLMs. This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by removing the separate vision encoder or tokenizer. For example, Fuyu-8B [6] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM, which markedly reduce inference latency. Inspired by this success, subsequent works [23, 64, 24, 47, 91, 82], further advance decode-only MLLMs by targeting their training processes and architectures. EvE [23] and VoRA [91] aligns the LLM’s hidden states with semantic features from a pre-trained ViT. However, directly using a smaller model (e.g., a ViT with hundreds of millions of parameters) as the teacher to distill knowledge into a significantly larger LLM (with several billion parameters) may restrict the expressive capacity of the LLM. Differently, Mono-InternVL [64] and EvEv2.0 [24] adopt a Mixture-of-Experts (MoE) framework, introducing a dedicated visual expert to handle visual-specific features more effectively. HoLVE [87] prepends a causal transformer to the LLM to explicitly convert both visual and textual inputs into a shared space. Despite these promising advancements, the overall training efficiency of decoder-only MLLMs remains a significant challenge. More importantly, the potential for the decoder-only architecture to create unified models that can seamlessly integrate multimodal understanding, generation, and even image editing capabilities remains a largely unexplored research avenue.
@@ 2.3 Unified Visual Understanding and Generation
Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier [96, 15, 102, 119, 94, 103, 50, 73, 11, 57, 21, 55]. Pioneering unified MLLMs such as Chameleon [88],Transfusion [119], emu3 [94], show-o [103] and Synergen-VL [50] utilize vison tokenizer (e.g., VQ-VAE) to convert images into discrete tokens, enabling seamless multimodal understanding and generation within a single model. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents. Janus series [96, 15] decouples visual encoding for understanding and generation using two separate encoders, but may compromise performance due to task conflicts in shared LLM parameter space. Metaqueries [73], BLIP3-O [11], Uniworld-V [57] assembles off-the-shelf specialized MLMMs and diffusion models by tuning adapters and learnable query tokens, which sacrifices true architectural unification for modularity. BAGEL [21] and Mogao [55] employ a Mixture-of-Transformers (MoT) architecture, dedicating different components to autoregressive text generation and diffusion-based visual generation. However, while powerful, this hybrid approach inherits the significant inference latency of diffusion models and still requires separate encoders and tokenizers during the inference.
@@ 2.4 Next Scale Prediction for Visual Generation
In contrast, our proposed OneCAT seamlessly unifies next-token prediction for text generation and next-scale prediction for visual generation within a single decoder-only transformer of the LLM, and proposes the scale-aware adapter to further exploit the scale-specific representation for enhanced visual generation.
@@ 3 OneCAT
Figure 3: Inference pipeline of OneCAT, a decoder-only autoregressive unified model that seamlessly supports multimodal understanding, image generation and image editing.
As depicted in Fig. 3, OneCAT employs a pure decoder-only architecture and eliminates the need for additional vision encoder or tokenizer during inference. This streamlined design significantly simplifies the model structure and reduces computational cost. Unlike existing unified MLLMs [15, 21] using semantic encoders like ViTs for understanding, we follow recent encoder-free MLLMs [64] and use a lightweight Patch Embedding layer to convert raw images into continuous visual tokens for efficient and lossless processing. The same Patch Embedding layer is also used to process reference images for editing tasks, thereby superseding separate VAE tokenizer used [21, 55] and further enhancing inference efficiency.
For text generation, OneCAT adheres to the Next-Token Prediction paradigm. For visual generation, it innovatively employs the Next-Scale Prediction [89], where images are generated in a hierarchical coarse-to-fine manner to progressively from the lowest to the highest resolution scales. Refer to Appendix A for more preliminaries for next-scale prediction.
3.1 Modality-MoE
OneCAT proposes a Modality-MoE architecture with three specialized feed-forward network (FFN) experts: a Text. FFN designed for text tokens for language comprehension, a Und. FFN designed for continuous visual tokens for visual understanding, and a Gen. FFN for discrete visual tokens for image generation. We employ a hard routing mechanism that assigns tokens to a specific expert based on their modality and the task at hand. Other QKV and attention layers are shared to promote parameter efficiency and robust cross-modal alignment for instruction-following.
OneCAT is initialized from the pre-trained Qwen2.5 LLM [105] to exploit its strong ability in language modeling. To construct the proposed Modality-MoE, we replicate the FFN layer from each Qwen2.5 transformer block to form three distinct experts. The core functionality for each multimodal task is handled as below.
@@ 3.1 Modality-MoE
We employ a simple yet effective patch embedding layer to convert raw images into continuous visual tokens. This layer consists of a $\times14$ convolution for image patchifying, a $\times2$ pixel unshuffle layer, and a two-layer Multilayer Perceptron (MLP) for projecting the visual features to match the LLM’s hidden state dimension.
@@
We leverage a pre-trained multi-scale VAE model from Infinity [31] to map images between the pixel and latent spaces. This VAE operates with a downsampling ratio of 16 and a latent channel size of 32, and uses a bitwise quantizer [118] to enlarge the vocabulary. During training, the VAE tokenizer encodes target images into multi-scale discrete visual tokens to serve as ground-truth. During inference, this tokenizer is not required, and only the detokenizer is used to reconstruct the image from generated multi-scale visual tokens.
@@
OneCAT seamlessly supports image editing task by leveraging a reference image and instruction input. The reference image is processed by the patch embedding layer, and the resulting continuous visual tokens are also routed to the Und. FFN to serve as the visual condition. The patch embedding layer provides a near-lossless representation of the reference image, and allows the LLM’s shallower layers to obtain low-level features for pixel-level consistency, while the deeper layers to extract high-level features for semantical comprehension. Guided by this hierarchical visual context, LLM predicts new discrete visual tokens autoregressively. This design enables powerful conditional generation without any architectural modifications, showing the versatility of our unified decoder-only design.
@@ 3.2 Scale-Aware Adapter for Hierarchical Generation
The tokens produced by the multi-scale VAE are inherently hierarchical (as shown in Appendix D). Specifically, lower-scale tokens mainly encode low-frequency global information such as color, illumination, and coarse structure, while higher-scale tokens capture high-frequency details including fine textures and intricate patterns. Processing these functionally divergent tokens equally with a shared Gen. FFN layer is not optimal and could limit the representation capacity.
To address this, we introduce the Scale-Aware Adapter (SAA), a novel architectural component integrated with the Gen. FFN. The SAA comprises a set of parallel modules that serve as skip connections over each linear layer of the Gen. FFN. Each SAA is dedicated to processing tokens from a specific scale, with the total number of SAA of a Gen. FFN matching the number of VAE scales. During inference, discrete visual tokens are routed to corresponding scale-specific adapters based on scale indices. To ensure parameter efficiency, each adapter is achieved using a low-rank decomposition (rank rr= 64), inspired by LoRA [33]. However, unlike LoRA that is typically used for fine-tuning, the SAA modules are jointly trained in an end-to-end manner as permanent components of the LLM.